OpenStudy (laurenarg):

What is the equation of the line that passes through (-2, 2) and (1, -4)? y = 2x - 2 y = -2x - 4 y = -2x - 2 y = 2x + 4

OpenStudy (blank):

explanation or answer?

OpenStudy (blank):

well first you should find the slope which can be found with formula \[\frac{ y_2-y_1 }{ x_2-x_1 } \]. The slope is -2. then replace this with the "a" in the formula y=ax+b. The next step is to find the "b" in the formula. To find this, just plug in one of the points in the problem into the equation. The answer is y=-2x-2.

OpenStudy (laurenarg):

Sorry... didn't see your answer. Thank you so much though..

OpenStudy (laurenarg):

Could you help with some more??

OpenStudy (laurenarg):

A circle has a diameter with endpoints (-8, 2) and (-2, 6). What is the equation of the circle? r2 = (x - 3)2 + (y + 4)2 r2 = (x - 5)2 + (y + 4)2 r2 = (x + 5)2 + (y - 4)2 r2 = (x + 3)2 + (y - 4)2